# Large Language Models for Toxic Language Detection in Low-Resource Balkan Languages

### v1

@@ Abstract
Online toxic language causes real harm, especially in regions with limited moderation tools. In this study, we evaluate how large language models handle toxic comments in Serbian, Croatian, and Bosnian, languages with limited labeled data. We built and manually labeled a dataset of 4,500 YouTube and TikTok comments drawn from videos across diverse categories, including music, politics, sports, modeling, influencer content, discussions of sexism, and general topics. Four models (GPT-3.5 Turbo, GPT-4.1, Gemini 1.5 Pro, and Claude 3 Opus) were tested in two modes: zero-shot and context-augmented. We measured precision, recall, F1 score, accuracy and false positive rates. Including a short context snippet raised recall by about 0.12 on average and improved F1 score by up to 0.10, though it sometimes increased false positives. The best balance came from Gemini in context-augmented mode, reaching an F1 score of 0.82 and accuracy of 0.82, while zero-shot GPT-4.1 led on precision and had the lowest false alarms. We show how adding minimal context can improve toxic language detection in low-resource settings and suggest practical strategies such as improved prompt design and threshold calibration. These results show that prompt design alone can yield meaningful gains in toxicity detection for underserved Balkan language communities.
@@
Toxic Language, Large Language Models, Natural Language Processing, Artificial Intelligence
@@ I-A Background & Problem Statement
Toxic language remains one of the most pressing challenges for social platforms, with significant psychological and societal impacts, particularly in politically sensitive or post-conflict regions [2]. The problem is amplified in online spaces due to anonymity, lack of accountability, and the speed at which harmful content can spread. These factors create environments where users feel emboldened to express hate or harassment they might avoid in offline settings [3]. While tech companies have invested in automated moderation systems, these systems often perform well only in high-resource languages like English [4]. For languages spoken across the Western Balkans, especially Serbian, Bosnian, and Croatian, moderation tools are less reliable, allowing toxic language and harassment to spread with limited oversight [5, 6]. This issue also extends to other closely related regional languages, which share similar linguistic and cultural characteristics but face the same lack of support in AI systems.
@@
This gap raises an important question about whether simple improvements, such as adding brief contextual information, could make these models more effective in practice. This paper examines how the performance of large language models changes when detecting toxic language in Balkan language comments, first using zero-shot prompts and then by adding minimal context. We test whether a short description of the video a comment responds to can help the model better interpret intent and tone. Our focus is on Serbian, Croatian, and Bosnian, where labeled datasets are limited and language variation presents real challenges for automated moderation.
We hypothesize that even a small amount of context can help models distinguish between ordinary criticism and toxic remarks, especially when comments involve cultural references, sarcasm, or platform-specific slang.
@@ I-B Objectives and Contributions
  • We introduce a manually labeled dataset of 4 500 YouTube and TikTok comments (1 500 per language), spanning a range of categories including music, politics, sports, modeling, influencers, sexism, and general topics.
  • We evaluate multiple versions of state of the art language models (GPT-3.5 Turbo, GPT-4.1, Gemini 1.5 Pro and Claude 3 Opus), comparing their performance with and without added context, and analyzing how contextual prompts affect false negatives across languages.
  • We provide qualitative insights into common failure cases, including misinterpretation of slang, sarcasm, and mixed script comments, which highlight the limits of zero-shot moderation in low resource settings.
We will share the labeled dataset upon request to enable reproducibility and facilitate follow-up studies.
@@ II-A Toxic Language Detection
Toxic language refers to a broad range of harmful expressions that may not meet the legal threshold of hate speech but can still cause significant harm, particularly in anonymous or large-scale online communication settings [1, 14]. A variety of machine learning classifiers such as Support Vector Machines (SVM), Random Forests, and deep learning architectures like Convolutional Neural Networks (CNNs) and Long Short-Term Memory networks (LSTMs) have been successfully applied to toxic language detection tasks [15].
@@
II-B LLM-based Methods
@@
Innovations in prompting strategies, such as chain of thought (CoT) prompting, further enhance LLM capabilities by guiding models to reason step by step, which has shown strong improvements in reasoning tasks [22]. Recent work has demonstrated that CoT-based frameworks, such as HateGuard, can significantly improve the detection of emerging and contextually complex hate speech in zero-shot settings [23]. However, despite this progress, the broader effectiveness of CoT prompting in detecting nuanced toxic language, particularly in cases that involve cultural references, sarcasm, or code switching, remains insufficiently studied. Existing research highlights these limitations. For example, LLMs still struggle with covert toxicity, including irony and coded insults, which reveals ongoing gaps in reliably identifying subtle forms of harm [24]. We do not experiment with CoT in this paper, but this suggests an important direction for future research.
II-C Low-Resource Languages and NLP
Processing natural language effectively in low-resource languages presents numerous challenges, notably due to the scarcity of extensive annotated datasets necessary for training robust machine learning models. These challenges are further amplified by the complex nature of many low-resource languages, which often include rich grammar, regional dialects, multiple writing systems, and frequent use of slang and idiomatic expressions [25]. In the case of Balkan languages such as Serbian, Bosnian, and Croatian, the combination of morphological complexity, dialectal diversity, and script variation (e.g., Serbian’s use of both Cyrillic and Latin) introduces additional hurdles for NLP pipelines [26, 27].
@@ II-C Low-Resource Languages and NLP
The current research addresses these identified gaps by systematically evaluating and comparing multiple state of the art LLMs specifically on toxic language detection tasks in Serbian, Bosnian, and Croatian, investigating both zero-shot performance and context-augmented classification to assess how well they function in practical toxic language detection tasks for low-resource Balkan languages.
@@ III-C Data Preparation
To ensure quality and privacy, all comments were preprocessed before model evaluation. Only comments written in Serbian, Bosnian, or Croatian were included. Comments consisting solely of emojis or entirely in non-target languages were excluded to maintain linguistic relevance. Preprocessing involved UTF-8 decoding, trimming unnecessary whitespace, and preserving emojis and punctuation, as these often convey important cues in online discourse. Personally identifiable information such as user IDs, comment IDs, and video IDs was not shared with the models. Mentions of public figures were retained during inference to preserve contextual meaning but were anonymized afterward to allow ethical data sharing.
@@ III-D Experimental Design
Each model was evaluated in two distinct prompting configurations: zero-shot and context-augmented. In the zero-shot scenario, the prompt presented the model with a user comment to determine toxicity:
@@ III-E Evaluation Metrics
To assess model performance in classifying toxic versus non-toxic comments, we used standard classification metrics: precision, recall, F1 score, accuracy and false positive rates. These metrics are commonly applied in NLP classification tasks to capture different aspects of prediction quality. Our dataset contained approximately 51% toxic and 49% non-toxic comments, allowing for a reliable application of this metric without requiring additional adjustments for class imbalance. All metrics were computed across the full dataset without separation by language, as the study treats Serbian, Bosnian, and Croatian jointly due to their high linguistic similarity. Additionally, we calculated the F1 score separately for each language to assess individual language performance. The computation of these metrics is step 9 in Figure 1. In the following section, we present and compare the results of all models under both zero-shot and context-augmented setups.
@@ IV-A Overview of Model Performance
TABLE I: Model performance on toxicity detection in low-resource languages, by configuration.
@@ IV-E Cost Comparison
We estimated API costs for 500 requests and then scaled to all 4 500 comments in our dataset (9 batches of 500), under both zero-shot and context-augmented modes. Figures assume current list prices at the time of writing and a temperature setting of 0. Table II shows per-request costs and scaled totals.
@@ IV-F Summary of Findings
Context-augmentation consistently boosted recall and F1 scores by roughly 0.07 to 0.10, although it sometimes led to higher false positive rates. The most balanced performer is Gemini 1.5 Pro context-augmented (GM-C), which achieved both the highest F1 score of 0.819 and the top accuracy of 0.823. GPT-4.1 zero-shot (G4-Z) remains the precision leader, making it suitable when false alarms carry a high penalty.
These results underscore that the optimal model configuration depends on the specific requirements of a moderation task—whether that’s minimizing false positives, maximizing detection of toxicity, or finding a middle ground. In the next section, we discuss how calibration, threshold tuning, and simple ensembling can further refine these trade-offs.
@@ V-A Key Findings
(“Why did they choose a sick person?”)
@@ V-B Practical Implications
    refers to a foul that prevented a counter-attack which could have led to a goal and secured a win for Serbia—yet GPT-3.5 Turbo context-augmented and Gemini 1.5 Pro context-augmented both misclassified it as toxic after seeing that background.
  • Threshold calibration and ensembling. Combining a high-precision zero-shot model (e.g., GPT-4.1 zero-shot) with a high-recall context-augmented model (e.g., GPT-3.5 Turbo context-augmented) can recover even more toxic content without unduly raising false positives.
@@ V-C Limitations
  • Static context. We provided only a short, manually written description of the video. For example, in one music video several comments focused on the singer’s hair—an angle not covered by our context prompt—so the models missed or mis-flagged them. Automated context extraction—such as summarizing user profiles, conversation threads, or external news links—might help capture those unanticipated topics and further improve performance.
@@ V-D Future Work
In summary, our discussion underscores that context-augmentation is a low-barrier, high-impact technique for improving toxicity detection in under-served languages. With further tuning and system integration, platforms can deploy these simple prompting strategies to make online spaces safer and more inclusive for Balkan language communities.
@@ VI Conclusion
  • Real-world scale. For GPT-3.5 Turbo alone, context-augmentation cut false negatives by 331 comments—hundreds of toxic remarks that would otherwise go unnoticed.
  • Edge cases highlight limits. Manually written context sometimes missed unexpected comment themes (e.g., singer hair or rebuttals of a sports foul), leading to over-flagging. Automating or broadening context extraction could address these gaps.
These findings demonstrate that simple prompt engineering—without any fine-tuning—can deliver state-of-the-art gains in toxicity detection for under-served languages. Looking forward, integrating dynamic context generation, exploring few-shot and chain-of-thought prompting, and incorporating lightweight fine-tuning or human-in-the-loop mechanisms promise further improvements. We hope this work encourages practitioners and researchers to adopt context-aware moderation pipelines and to extend these techniques to other low-resource and culturally rich language settings.

### v2

@@ Abstract
Online toxic language causes real harm, especially in regions with limited moderation tools. In this study, we evaluate how large language models handle toxic comments in Serbian, Croatian, and Bosnian, languages with limited labeled data. We built and manually labeled a dataset of 4,500 YouTube and TikTok comments drawn from videos across diverse categories, including music, politics, sports, modeling, influencer content, discussions of sexism, and general topics. We demonstrate how adding minimal context can improve toxic language detection in low-resource settings and suggest practical strategies such as improved prompt design and threshold calibration. Four models (GPT-3.5 Turbo, GPT-4.1, Gemini 1.5 Pro, and Claude 3 Opus) were tested in two modes: zero-shot and context-augmented. We measured precision, recall, F1 score, accuracy and false positive rates. Including a short context snippet raised recall by about 0.12 on average and improved F1 score by up to 0.10, though it sometimes increased false positives. The best balance came from Gemini in context-augmented mode, reaching an F1 score of 0.82 and accuracy of 0.82, while zero-shot GPT-4.1 led on precision and had the lowest false alarms. These results show that prompt design alone can yield meaningful gains in toxicity detection for underserved Balkan language communities.
@@
Large Language Models, Natural Language Processing, Artificial Intelligence, Toxic Language
@@ I-A Background & Problem Statement
Toxic language remains one of the most pressing challenges for social platforms, with significant psychological and societal impacts, particularly in politically sensitive or post-conflict regions [2]. The problem is amplified in online spaces due to anonymity, lack of accountability, and the speed at which harmful content can spread. These factors create environments where users feel emboldened to express hate or harassment they might avoid in offline settings [3]. While tech companies have invested in automated moderation systems, these systems often perform well only in high-resource languages like English [4]. Most rely on keyword filters or classifiers trained on large datasets, but they fall short in low-resource languages due to limited data, cultural nuance, and evolving online language patterns [5]. For languages spoken across the Western Balkans, especially Serbian, Bosnian, and Croatian, moderation tools are less reliable, allowing toxic language and harassment to spread with limited oversight [6].
@@
This gap raises an important question about whether simple improvements, such as adding brief contextual information, could make these models more effective in practice. This paper examines how the performance of large language models changes when detecting toxic language in Balkan language comments, first using zero-shot prompts and then by adding minimal context. We test whether a short description of the video a comment responds to can help the model better interpret intent and tone. Our hypothesis is that even a small amount of context can help models distinguish between ordinary criticism and toxic remarks, especially when comments involve cultural references, sarcasm, or platform-specific slang.
To our knowledge, no prior study has systematically evaluated the performance of state-of-the-art large language models for toxic language detection in Serbian, Croatian, and Bosnian, making this a notable gap in the current research landscape.
@@ I-B Objectives and Contributions
  • Introducing a manually labeled dataset of 4 500 YouTube and TikTok comments (1 500 per language), spanning a range of categories including music, politics, sports, modeling, influencers, sexism, and general topics.
  • Evaluating multiple versions of state-of-the-art language models (GPT-3.5 Turbo, GPT-4.1, Gemini 1.5 Pro and Claude 3 Opus), comparing their performance with and without added context, and analyzing how contextual prompts affect false negatives across languages.
  • Providing qualitative insights into common failure cases, including misinterpretation of slang, sarcasm, and mixed script comments, which highlight the limits of zero-shot moderation in low resource settings.
@@ II-A Toxic Language Detection
Prior work on toxic language detection has focused on various machine learning approaches and challenges, especially in high-resource languages. A variety of classifiers such as Support Vector Machines (SVM), Random Forests, and deep learning architectures like Convolutional Neural Networks (CNNs) and Long Short-Term Memory networks (LSTMs) have been successfully applied to this problem [14, 15].
@@
II-B Large Language Model-based Methods
@@
Innovations in prompting strategies, such as chain-of-thought (CoT) prompting, further enhance LLM capabilities by guiding models to reason step by step, which has shown strong improvements in reasoning tasks [22]. Recent work has demonstrated that CoT-based frameworks, such as HateGuard, can significantly improve the detection of emerging and contextually complex hate speech in zero-shot settings [23]. However, despite this progress, the broader effectiveness of CoT prompting in detecting nuanced toxic language, particularly in cases that involve cultural references, sarcasm, or code switching, remains insufficiently studied. Existing research highlights these limitations. For example, LLMs still struggle with covert toxicity, including irony and coded insults, which reveals ongoing gaps in reliably identifying subtle forms of harm [24]. We do not experiment with CoT in this paper, but this suggests an important direction for future research.
II-C Low-Resource Languages and Natural Language Processing
Processing natural language effectively in low-resource languages presents numerous challenges, notably due to the scarcity of extensive annotated datasets necessary for training robust machine learning models. These challenges are further amplified by the complex nature of many low-resource languages, which often include rich grammar, regional dialects, multiple writing systems, and frequent use of slang and idiomatic expressions [25]. In the case of Balkan languages such as Serbian, Bosnian, and Croatian, the combination of morphological complexity, dialectal diversity, and script variation (e.g., Serbian’s use of both Cyrillic and Latin) introduces additional hurdles for Natural Language Processing (NLP) pipelines [26, 27].
@@ II-C Low-Resource Languages and Natural Language Processing
This study expands on the above findings to empirically evaluate large language models on toxic comment detection in Balkan languages.
@@ III-C Data Preparation
To ensure quality and privacy, all comments were preprocessed before model evaluation. Comments consisting solely of emojis or entirely in non-target languages were excluded to maintain linguistic relevance. Preprocessing involved UTF-8 decoding, trimming unnecessary whitespace, and preserving emojis and punctuation, as these often convey important cues in online discourse. Personally identifiable information such as user IDs, comment IDs, and video IDs was not shared with the models. Mentions of public figures were retained during inference to preserve contextual meaning but were anonymized afterward to allow ethical data sharing.
@@ III-D Experimental Design
We compared model behavior across two prompting strategies: zero-shot (where the model receives only the comment and a task description) and context-augmented (where the model also receives a short description of the video to help interpret the comment). In the zero-shot scenario, the prompt presented the model with a user comment to determine toxicity:
@@ III-E Evaluation Metrics
To assess model performance in classifying toxic versus non-toxic comments, we used standard classification metrics: precision, recall, F1 score, accuracy and false positive rate (FPR). These metrics are commonly applied in NLP classification tasks to capture different aspects of prediction quality. Our dataset contained approximately 51% toxic and 49% non-toxic comments, allowing for a reliable application of this metric without requiring additional adjustments for class imbalance. We computed all metrics across the full dataset, treating Serbian, Bosnian, and Croatian jointly due to their high linguistic similarity. In addition, we reported F1 scores separately for each language to highlight any performance differences between them. The computation of these metrics is step 9 in Figure 1. In the following section, we present and compare the results of all models under both zero-shot and context-augmented setups.
@@ IV-A Overview of Model Performance
TABLE I: Model Performance on Toxicity Detection in Low-Resource Languages by Configuration
@@ IV-E Cost Comparison
We measured API costs for 500 requests and then scaled to all 4 500 comments in our dataset (9 batches of 500), under both zero-shot and context-augmented modes. Figures assume list prices at the time of writing and a temperature setting of 0. Table II shows total API costs for 500 and 4 500 requests under each configuration.
@@ IV-F Summary of Findings
Context augmentation consistently boosted recall and F1 scores by roughly 0.07 to 0.10, although it sometimes led to higher false positive rates. The most balanced performer is Gemini 1.5 Pro context-augmented (GM-C), which achieved both the highest F1 score of 0.819 and the top accuracy of 0.823. GPT-4.1 zero-shot (G4-Z) remains the precision leader, making it suitable when false alarms carry a high penalty.
These results underscore that the optimal model configuration depends on the specific requirements of a moderation task, whether that’s minimizing false positives, maximizing detection of toxicity, or finding a middle ground. In the next section, we discuss how calibration, threshold tuning, and simple ensemble approach can further refine these trade-offs.
@@ V-A Key Findings
(“Why did they elect a sick person?”)
@@ V-B Practical Implications
    was made in response to a football match video, referring to a foul that stopped a counter-attack which could have led to a goal and secured a win for Serbia. Yet, GPT-3.5 Turbo context-augmented and Gemini 1.5 Pro context-augmented both misclassified it as toxic after seeing that background.
  • Threshold calibration and ensemble approach. Combining a high-precision zero-shot model (e.g., GPT-4.1 zero-shot) with a high-recall context-augmented model (e.g., GPT-3.5 Turbo context-augmented) can recover even more toxic content without unduly raising false positives.
@@ V-C Limitations
  • Static context. We provided only a short, manually written description of the video. For example, in one video featuring a well-known TikToker who was throwing money on the street, several comments unexpectedly focused on his hairstyle. This detail was not included in our context prompt, as it seemed irrelevant at the time. As a result, the models missed or misclassified some of these comments. Automated context extraction, such as summarizing user profiles, conversation threads, or external news links might help capture those unanticipated topics and further improve performance.
@@ V-D Future Work
In summary, our discussion underscores that context augmentation is a low-barrier, high-impact technique for improving toxicity detection in under-served languages. With further tuning and system integration, platforms can deploy these simple prompting strategies to make online spaces safer and more inclusive for Balkan language communities.
@@ VI Conclusion
  • Real-world scale. For GPT-3.5 Turbo alone, context-augmentation cut false negatives by 331 comments, hundreds of toxic remarks that would otherwise go unnoticed.
  • Edge cases highlight limits. Manually written context sometimes missed unexpected comment themes (e.g. rebuttals of a sports foul), leading to over-flagging. Automating or broadening context extraction could address these gaps.
These findings demonstrate that simple prompt engineering, without any fine-tuning, can deliver state-of-the-art gains in toxicity detection for under-served languages. Looking forward, integrating dynamic context generation, exploring few-shot and chain-of-thought prompting, and incorporating lightweight fine-tuning or human-in-the-loop mechanisms promise further improvements. We hope this work encourages practitioners and researchers to adopt context-aware moderation pipelines and to extend these techniques to other low-resource and culturally rich language settings.
Conflict of Interest
The authors declare no conflict of interest.
Data Availability